In a study of long bones in a variety of mammalian species, a regression analysis of y = humerus bone circumference (cm) on xWhat is the upper limit for the Cl?
Answer #1
x0             = 40
predcited value at X=40: 38.94
std error prediction interval= s*√(1+1/n+(x0-x̅)2/Sxx) = 6.7312
for 95 % CI value of t= 1.9730
margin of error E=t*std error                            = 13.28
lower prediction bound=sample mean-margin of error = 25.6594
Upper prediction bound=sample mean+margin of error= 52.2206

lower limit of the CI =25.6594

upper limit of the CI = 52.2206